2x + 4y =10
3x - 2y = -9


Answer by checkley77(12844)

2x + 4y =10
3x - 2y = -9 multiply by 2 & add.
-----
2x+4y=10
6x-4y=-18
-------------------------------
8x=-8
x=-8/8
x=-1 ans.
2*-1+4y=10
-2+4y=10
4y=10+2
4y=12
y=12/4
y=3 ans.
Proof:
3*-1-2*3=-9
-3-6=-9
-9=-9